# Thermodynamics of the dipole-octupole pyrochlore magnet CeHfO in applied magnetic fields

### v1

@@ V Thermodynamics and effects of misalignment for magnetic fields applied along the [111] direction
At low fields and low temperatures we expect the presence of a putative U(1) QSL – the evidence being based on the lack of magnetic ordering, e.g., the absence of any Bragg peaks in neutron scattering, and the observation of a continuum of spin excitations instead. On heating, such a state simply evolves into the disordered paramagnetic state by a crossover, explaining why there is only a single bump in the specific heat on a scale that depends both on the ’s and .
At higher fields, the [111] magnetic field, selects a unique symmetry broken state ‘3-in–1-out’ (‘1-in–3-out’) state. It too is related to the disordered state at high temperature only via a crossover. However, if one looks closely at the data, two fused features appear in the specific heat at high fields - this can be attributed to the coupling of the field to sublattice 0 being inequivalent to that of sublattices 1,2,3.
We now investigate the effect of misalignment of [111] fields. We parameterize the field components for the misalignment using Eq. (IV.3) where and represent the deviation angles defined previously in Fig. 1. Here, the case of and represents perfect alignment along the [111] direction.
The role of misalignment, while still important in controlling the height of the specific heat bumps, in general, is greatly diminished compared to the case of the [110] field. We have fitted the misalignment angles and for the case of [111] specific heat using a fitting procedure similar to the case of [110] field. In Figure 10 we plot the specific heat profiles computed within FTLM using these optimum values of misalignment angles. Even though the fitting procedure results in a misalignment of about 10 degrees, there is very little to distinguish the results with those for the case of no misalignment (0 degrees) i.e. the model does not improve significantly the global agreement with the data. This observation is consistent with torque calculations that indicate that a field applied close to the [111] direction is expected to realign this sample direction along the field, as illustrated in Fig. 7.
@@ VI Conclusions
Dipole-octupole systems serve as a rich source of pseudospin-1/2 physics and offer access to phases where quantum fluctations are enhanced by geometric frustration, an amalgamation of conditions that are conducive for the occurrence of quantum spin liquidity [46, 59, 62, 63]. In contrast to the familiar dipolar systems, the relative inertness of the octupolar components to applied magnetic fields has profound consequences, and this paper explored a set of unusual effects emerging from such physics. While the current paper focused primarily on unpublished experimental data for CeHfO, we believe that most conclusions can also transfer to its sister compound CeZrO [53, 54, 56].
Our experiments on CeHfO(neutron diffraction, specific heat measured with two independent setups and magnetization) and corresponding numerical simulations agree remarkably well with each other in the low field limit, providing confidence in the microscopic description and our previously determined parameters [57]. However, data in the high field limit, which in principle should be the easiest to explain, pose puzzles which we have addressed here. In particular, our experimental data for the specific heat of CeHfO exhibit two distinct bumps indicating a two-step release of entropy. While it may be tempting to interpret the low temperature bump as a precursor of a phase transition, our current numerical analysis suggests that the observed behavior is primarily due to Zeeman coupling of the field to chains which is otherwise forbidden in idealistic settings. A more refined look at the experiments and theory was required to decipher the origins of our experimental observations.
We have explored three scenarios which would make the chains directly visible to the applied [110] field. The first explores the effective interaction between the chains mediated by the spins locked to the [110] field, which emerges from an order-by-disorder mechanism proposed by Ref. [68]. We deduce that the theory would predict a low temperature specific heat feature that decreases with field strength, rather than the increase seen in our experiments. Second, we investigated the possible importance of symmetry-allowed terms beyond the linear Zeeman term, in principle, a cubic Zeeman term directly couples spins to the field. Using CEF and spin-orbit coupled ab initio calculations we find that this effect is too small.
@@ VII Acknowledgments
We thank O. Tchernyshyov, R. Moessner, B. Gaulin, Y. Kim, S. Zhang and A. Patri for insightful discussions and correspondence. We also thank A. Scheie for help with the PyCEF software. H.J.C. was supported by NSF CAREER grant No. DMR-2046570. A.B. and H.J.C. thank Florida State University and the National High Magnetic Field Laboratory for support. The National High Magnetic Field Laboratory is supported by the National Science Foundation through NSF/DMR-2128556 and the state of Florida. We also thank the Research Computing Center (RCC) and Planck cluster at Florida State University for computing resources. H.Y. and A.H.N. were supported by the National Science Foundation Division of Materials Research under the Award DMR-1917511. N.G. and J.A.Q. acknowledge the support of the Canada First Research Excellence Fund (CFREF). We acknowledge funding from the Swiss National Science Foundation (project No. 200021_179150) and European Commission under grant agreement no. 824109 European ‘Microkelvin Platform’. This work is also based on experiments performed at the Swiss spallation neutron source SINQ (Paul Scherrer Institute, Switzerland)
Appendix A
@@
With the CEF parameters mentioned in the main text, the specific heat for a single tetrahedron in the presence of a [110] field is plotted in Fig. 11 by using Eq. 10. It is clear that the ab-initio Hamiltonian is able to capture the high-temperature part of the specific heat accurately, corroborating our parameter set. It should be noted that at smaller values of field ( T) the Zeeman term in the ab-intio Hamiltonian for sub-lattice 1 and 2 are essentially zero and, therefore, the two sites do not contribute to the specific heat. Hence, the spins alone explain the field-dependent behaviour of the specific heat at higher temperatures. This is also clearly realised by noting that the evaluated entropy (per site) for the single tetrahedron as a function of temperature with the field along [110] direction and at very high temperature is instead of (see Fig. 11b).
@@
We confirm the finding of the weak coupling of the sites, by calculating . To do so we evaluate the energy splitting of the lowest lying doublet in the presence of [110] field using Eq. 10. By fitting to an appropriately defined cost function, this energy difference curve directly gives us , which we find to be meV.
@@
In Fig. 11 we show the energy splitting for sublattice 2 obtained by using the ab-initio Hamiltonian and from the effective theory with meV. Even at 50 Tesla the energy difference is of the order of 0.01 meV, suggesting that the system is still in the perturbative regime, and that the estimate of meV (stated in the main text) from directly fitting the low field data is unrealistic.
Appendix B
@@
The specific heat data presented in this work, for the highest magnetic fields, were taken using the PPMS He specific heat setup. From our analysis, we argue that relatively large tilts of the sample platform can be induced by the torque of the sample when applying magnetic field in cryogenic conditions. As an example of this we show in Fig. 12 the PPMS He specific heat platform after an unrelated experiment performed using a similar setup on a sample of composition HoGa. The sample was cooled down without magnetic field, after which a moderate magnetic field was applied for measurements, then removed before warming up the sample. The tilt of the sample platform is approximately 30 degrees. In such an extreme case, the sample torque triggers changes to the platform going beyond the ‘plastic limit’ of the assembly, which thus can be observed after taking the setup out of the cryomagnet. The value of deduced from the present experiments on CeHfO remain smaller than what is exemplified in Fig. 12.
Appendix C
@@
In order to capture the behaviour of the magnetization as a function of a magnetic field applied along the [110] direction, we also considered the effect of the non-magnetic Ce impurities. To do so, we used the two CEF models developed in [55]: one where the Ce ion experiences an ideal D environment and a second one where the D symmetry is broken. These two models were subsequently fitted to the INS and magnetic susceptibility data [55]. A notable difference between the two models is the reduced Ising anisotropy of the defective model, which translates into a g-tensor containing more terms than . For both models, we compute the magnetization for a field applied along the [110] direction. Consequently to the reduced anisotropy, the defective model shows a higher signal compared to the perfect one.
@@
Assuming diluted defects, a system with sites has 0.25 number of sites of each sublattice. If the amount of impurity is 10% in the form of Ce and are homogeneously distributed then 0.025 number of sites on each sublattice are nonmagnetic Ce ions. Furthermore if these defects are far from each other and any magnetic site at most see only a single impurity, then 0.15/0.075 number of sites of each sublattice type experience a defective/perfect crystal field environment (this is because any single impurity at any sublattice will affect the environment of 6 neighboring cerium atoms). This means that 60% of Ce experience a defective environment, and the remaining 30% experience a perfect CEF environment.
@@
Consequently, summing the magnetization curves scaled by 0.3 and 0.6 for the perfect and defective models, respectively, should reproduce the experimental data. However this approach falls short of capturing the measured magnetization, as can be seen in Fig.13. As expected, the high-temperature data are reasonably well captured by this model, as exchange interactions are negligible in this regime. Conversely, the low temperature data are not well reproduced, particularly at low field where the exchange terms are dominant. Furthermore, not only is the overall magnitude of the computed signal is too small, but the slope observed at high magnetic fields (above 4 T) is also flatter than that of the experimental data.
@@
Based on these observations, we argue that although the presence of impurities affects the magnetization behaviour, its impact is minor and does not permit to accurately reproduce our experimental data.

### v2

@@ V Thermodynamics and effects of misalignment for magnetic fields applied along the [111] direction
At low fields and low temperatures we expect the presence of a putative U(1) QSL – the evidence being based on the lack of magnetic ordering, e.g., the absence of any Bragg peaks in neutron scattering, and the observation of a continuum of spin excitations instead [61]. On heating, such a state evolves into the disordered paramagnetic state by a crossover, potentially explaining why there is only a single bump in the specific heat on a scale that depends both on the ’s and .
At higher fields, the [111] magnetic field selects a unique symmetry broken state ‘3-in–1-out’ (‘1-in–3-out’) state. It too is related to the disordered state at high temperature only via a crossover. However, on looking closely at the data, we observe two fused features in the specific heat at high fields - this can be attributed to the coupling of the field to sublattice 0 being inequivalent to that of sublattices 1,2,3.
We now investigate the effect of misalignment of [111] fields. We parameterize the field components for the misalignment using Eq. (7) where and represent the deviation angles defined previously in Fig. 1. Here, the case of and represents perfect alignment along the [111] direction.
The role of misalignment, while still important in controlling the height of the specific heat peaks, in general, is greatly diminished compared to the case of the [110] field. We have fitted the misalignment angles and for the case of specific heat in a [111] field using a fitting procedure similar to the case of a [110] field. In Fig. 9 we plot the specific heat profiles, computed with FTLM, using these optimum values of misalignment angles. Even though the fitting procedure results in a misalignment of about 10 degrees, there is very little to distinguish the results with those for the case of no misalignment (0 degrees) i.e. the model does not improve significantly the global agreement with the data. This observation is consistent with torque calculations that indicate that a field applied close to the [111] direction is expected to realign this sample direction along the field, as illustrated in Fig. 6.
@@ VI Conclusions
Dipole-octupole systems serve as a rich source of pseudospin-1/2 physics and offer access to phases where quantum fluctuations are enhanced by geometric frustration, an amalgamation of conditions that are conducive for the occurrence of quantum spin liquidity [51, 64, 53, 52]. In contrast to the familiar dipolar systems, the relative inertness of the octupolar components to applied magnetic fields has profound consequences, and this paper explored a set of unusual effects emerging from such physics. While the current paper focused primarily on previously unpublished experimental data for Ce2Hf2O7, we believe that most conclusions should also transfer to its sister compound Ce2Zr2O7 [46, 47, 56].
Our experiments on Ce2Hf2O7 (neutron diffraction, specific heat measured with two independent setups and magnetization) and corresponding numerical simulations agree remarkably well with each other in the low field limit, providing confidence in the microscopic description and our previously determined parameters [61]. However, data in the high field limit, which in principle should be the easiest to explain, pose puzzles which we have addressed here. In particular, our experimental data for the specific heat of Ce2Hf2O7 exhibit two distinct peaks indicating a two-step release of entropy. The low temperature peak could be associated with a phase transition, as has been suggested recently [55, 54] and whose details remain to be investigated further– our current numerical analysis suggests that its observed movement with field is primarily due to Zeeman coupling of the field to chains which is otherwise forbidden in idealistic settings. A more refined look at the experiments and theory was required to decipher the origins of our experimental observations.
We have explored three scenarios which would make the chains directly visible to the applied [110] field. The first explores the effective interaction between the chains mediated by the spins locked to the [110] field, which emerges from an order-by-disorder mechanism proposed by Ref. [53]. We deduce that the theory would predict a low temperature specific heat feature that decreases with field strength, rather than the increase seen in our experiments. Second, we investigated the possible importance of symmetry-allowed terms beyond the linear Zeeman term, in principle, a cubic Zeeman term directly couples spins to the field. Using CEF and spin-orbit coupled ab-initio calculations we find that this effect is too small.
@@
We conclude by noting that our work benefited from a close synergy between experiment and theory, the latter employed both ab-initio and many-body calculations. Given that multiple competing scales exist in multipolar magnetic systems, careful comparisons are indispensable for determining the effective microscopic description [61] and also identifying the origins of the observed responses in applied magnetic fields. More generally, many of the strategies employed in this work should be transferable to other magnetic materials (in the dipole-octupole family and beyond), enabling a better qualitative and quantitative understanding of their physical properties.
@@ VII Acknowledgments
We thank O. Tchernyshyov, R. Moessner, B. Gaulin, Y. Kim, S. Zhang and A. Patri for insightful discussions and correspondence. We also thank A. Scheie for help with the PyCEF software [75]. H.J.C. was supported by National Science Foundation CAREER grant No. DMR-2046570. A.B. and H.J.C. thank Florida State University and the National High Magnetic Field Laboratory for support. The National High Magnetic Field Laboratory is supported by the National Science Foundation through NSF/DMR-2128556 and the state of Florida. We also thank the Research Computing Center (RCC) and Planck cluster at Florida State University for computing resources. H.Y. and A.H.N. were supported by the National Science Foundation Division of Materials Research under the Award DMR-1917511. The theoretical analysis by A.H.N. was supported by the U.S. Department of Energy, Office of Basic Energy Sciences under Award no. DE-SC0025047. N.G. and J.A.Q. acknowledge the support of the Canada First Research Excellence Fund (CFREF). We acknowledge funding from the Swiss National Science Foundation (project No. 200021_179150) and European Commission under grant agreement no. 824109 European ‘Microkelvin Platform’. This work is also based on experiments performed at the Swiss spallation neutron source SINQ (Paul Scherrer Institute, Switzerland)
Appendix A Revisiting alternative scenarios
A.1 Details of the first scenario: Indirect coupling of spins
We highlight the details of the framework structured by PMB who extensively studied the classical phase diagram of dipole-octupole pyrochlores in a [110] magnetic field for a wide range of parameter sets. Here we focus on the case where is dominant, since it is directly pertinent to the study of Ce2Zr2O7 and Ce2Hf2O7. It is worth noting that the argumentation would essentially remain similar if is dominant.
The subextensive degeneracy is split by coupling between the chains, each representing an Ising spin, which themselves form an anisotropic triangular lattice. This coupling is a result of finite but small while keeping the [110] field large. In this regime, corresponding to the so-called ‘chain Y’ phase, the calculated couplings between nearest neighbor Ising spins are given by Eq. 5. As mentioned in section IV.1, this functional dependence suggests that the order-disorder transition of the chains should occur at lower temperatures upon increasing the field strength, which is opposite to the trend we observe in experiments.
We take this opportunity to clarify that the PMB procedure completely ignores quantum mechanical fluctuations within a chain. If these fluctuations are important, it is not apparent that it is justified to treat an entire chain as a single Ising degree of freedom. At large [110] fields the spins are essentially locked and can be treated classically. However, sites on a single chain are interacting quantum mechanically with their immediate nearest neighbors on the same chain, and this intrinsically selects a state on the scale of the magnetic coupling strengths (which are at the scale of 0.4 K or less), essentially independent of the magnetic field strength. We see evidence for our assertion in Fig. 10 where we plot the location of the low temperature feature in the specific heat as a function of field strength. We see only a small increase with increasing for low , and essentially independence at large .
A.2 Details of the second scenario: Higher order Zeeman coupling
Here we discuss the details of the second scenario (mentioned in section IV.2) and the fitting procedure used to estimate in Eq. 6. As a first exercise we attempted to fit the effective Hamiltonian parameters, along with the additional term, to our specific heat data. We found to be somewhat large ( meV-2), and our fits could only reproduce the experimental features at low fields.
In order to assess the possible importance of higher-order field terms and extract the directly, we performed diagonalization of the single ion ab-initio Hamiltonian which is a sum of the spin-orbit coupling term, the crystal field Hamiltonian and the Zeeman term in the space spanned by and manifold (i.e., Hilbert space of size 14). We ignore interactions between the ions. For a single ion, the form of the crystal field Hamiltonian is
For Ce2Hf2O7, , , , , and in meV. are the Stevens operators [77, 78].
These values of crystal field parameters were determined by fitting the CEF Hamiltonian parameters to inelastic neutron spectroscopy and magnetic susceptibility data measured on a powder sample of Ce2Hf2O7 as reported in [60]. Although the fit was somewhat under-constrained, it showed good agreement with the experimental data and produced reasonable uncertainties on the CEF parameters, making them a suitable basis for our model. Furthermore, variations within these uncertainties are unlikely to significantly alter the wavefunction and the associated magnetic g-tensor.
The spin-orbit coupling part of the Hamiltonian is given by
Here, , which determines the energy scale of the spin-orbit interaction, was set at 80.3 meV.
The bare Zeeman term is given by
With the CEF parameters mentioned above, the specific heat for a single tetrahedron in the presence of a [110] field is plotted in Fig. 11(a) by using Eq. 11.
@@ A.2 Details of the second scenario: Higher order Zeeman coupling
The ab-initio Hamiltonian is able to capture the high-temperature part of the specific heat accurately. It should be noted that at smaller values of field ( T) the Zeeman terms in the ab-initio Hamiltonian for sublattice 1 and 2 ( spins) are essentially zero and, therefore, the two sites do not contribute to the specific heat. Hence, the spins alone explain the field-dependent behavior of the specific heat at higher temperatures. This is also clearly realized by noting that the evaluated entropy (per site) for the single tetrahedron as a function of temperature with the field along [110] direction and at very high temperature is instead of (see Fig. 11b).
@@
We confirm the finding of the weak coupling of the sites, by calculating . To do so, we evaluate the energy splitting of the lowest lying doublet in the presence of [110] field using Eq. 11. By fitting to an appropriately defined cost function, this energy difference curve directly gives us , which we find to be meV-2.
@@
In Fig. 11c we show the energy splitting for sublattice 2 obtained by using the ab-initio Hamiltonian and from the effective theory with meV-2. Even at 50 Tesla the energy splitting of the doublet is only of the order of 0.01 meV, suggesting that the system is still in the perturbative regime, and that the estimate of meV-2 from directly fitting the low field specific heat data is unrealistic.
Appendix B Specific heat measurements at high fields
@@
The specific heat data presented in this work, for the highest magnetic fields, were recorded using the PPMS 3He specific heat setup. From our analysis, we argue that relatively large tilts of the sample platform can be induced by the torque of the sample when applying magnetic field in cryogenic conditions. As an example of this we show in Fig. 12 the PPMS 3He specific heat platform after an unrelated experiment performed using a similar setup on a sample of composition HoGa. The sample was cooled down without magnetic field, after which a moderate magnetic field was applied for measurements, then removed before warming up the sample. The tilt of the sample platform is approximately 30 degrees. In such an extreme case, the sample torque triggers changes to the platform going beyond the ‘plastic limit’ of the assembly, which thus can be observed after taking the setup out of the cryomagnet. The value of deduced from the present experiments on Ce2Hf2O7 remain smaller than what is exemplified in Fig. 12.
Appendix C Impact of nonmagnetic impurities on magnetization
@@
In order to capture the behavior of the magnetization as a function of a magnetic field applied along the [110] direction, we also considered the effect of the non-magnetic Ce4+ impurities. To do so, we used the two CEF models developed in [60]: one where the Ce3+ ion experiences an ideal D3d environment and a second one where the D3d symmetry is broken. These two models were subsequently fitted to the INS and magnetic susceptibility data [60]. A notable difference between the two models is the reduced Ising anisotropy of the model with impurities, which translates into a -tensor containing more non-zero contributions beyond .
@@
A system with sites has 0.25 number of spins on each sublattice. We assume a dilute density of impurities (in the form of nonmagnetic Ce4+ in the ballpark of 10) that are homogeneously distributed - this means 0.025 sites on each sublattice are nonmagnetic. If these impurities are far from each other and any magnetic site at most sees only a single impurity, then 0.15 (0.075) sites of each sublattice type experience a “defective” (perfect) crystal field environment; this is because any single impurity on any sublattice will affect the environment of 6 neighboring cerium atoms. This means that 60% of Ce experience a defective environment, and the remaining 30% experience a perfect CEF environment.
@@
If this effect were dominant, we anticipate that summing the magnetization curves scaled by 0.3 and 0.6 for the pure and impurity models, respectively, would reproduce the experimental data. However, we find that this approach falls short of capturing the measured magnetization, as can be seen in Fig.13. As expected, the high-temperature data are reasonably well captured by this model, since exchange interactions are negligible in this regime. However, the low-temperature data are not well reproduced, particularly at low field strengths where the exchange terms are dominant. Additionally, at intermediate temperatures, not only is the magnitude of the computed high-field magnetization too small, but the computed slope (above 4 T) is also smaller than that observed in the experiment.
@@
Based on these observations, we argue that although the presence of impurities does affect the magnetization, its impact is minor and alone does not accurately reproduce our experimental data.